OpenStudy (anonymous):

If an object is pushed 5m along a certain incline, its height above the ground increases by 2m. If the incline were frictionless it would require a force of 196N exerted parallel to the incline to push a 50kg crate up the incline at a constant speed. How much work would be done in pushing the crate up the incline?

OpenStudy (anonymous):

Work is equal to force times displacement times the cosine of the angle. W=f*s*cos(theta) The tricky part of your problem is finding theta. Since I'm on my phone I can't draw the picture as I would like to, but you are going to take the arctan of 5/2 to find the angle. Tan^-1(5/2)=calculated angle. The rest of this problem is plug and chug. W=196N*5m*calculated angle

OpenStudy (anonymous):

I'm sorry, I said arc tan and it should be arcsin because sin=opp/hyp. And work will be: W=196N*5m*cos(calculated angle) Sorry for the mistakes. All should be correct now